### v2

# Benchmarking Pretrained Vision Embeddings for Near- and Duplicate Detection in Medical Images

Abstract
Near- and duplicate image detection is a critical concern in the field of medical imaging. Medical datasets often contain similar or duplicate images from various sources, which can lead to significant performance issues and evaluation biases, especially in machine learning tasks due to data leakage between training and testing subsets. In this paper, we present an approach for identifying near- and duplicate 3D medical images leveraging publicly available 2D computer vision embeddings. We assessed our approach by comparing embeddings extracted from two state-of-the-art self-supervised pretrained models and two different vector index structures for similarity retrieval. We generate an experimental benchmark based on the publicly available Medical Segmentation Decathlon dataset. The proposed method yields promising results for near- and duplicate image detection achieving a mean sensitivity and specificity of 0.9645 and 0.8559, respectively.
I Introduction
Near- and duplicate image detection aims to find similar or identical images in a large data corpus and has many applications in web image retrieval and forensic tasks. Most image (near-) duplicate detection algorithms rely on matching images based on distance measures of associated image embeddings. In order to classify (near-) duplicates and non-duplicates typically a distance threshold is used. The most common handcrafted embeddings are SIFT [1] and SURF [2] features. Recent works leverage Deep Learning features [3, 4, 5] or combine with handcrafted features such as SIFT [6] to construct features at local and global level.
In medical imaging, identical or slightly preprocessed images can be sold by different data brokers or even by the same vendor for different tasks. This is in particular risky when machine learning approaches are used. The model performance may be negatively impacted by (near-) duplicates or even evaluation biases might occur due to data leakage between train and test sets as observed in [7]. Developing (near-) duplicate detection algorithms for medical images comes with several challenges. First, most medical images are acquired in 3D and need to be projected to a low-dimensional representation for similarity search. While there are plenty of powerful pre-trained 2D image feature extractors available [8, 9], there is much less work related to 3D feature extraction. Naively matching slice-wise embeddings is not trivial. For example, border slices often contain uninformative background voxels which may cause mismatches with border slices from other cases regardless of anatomy or modality. This has direct implications for robustly calibrating the similarity threshold for duplicate identification. In our approach, we aim to leverage 2D feature extractors pretrained on natural images and propose a count-based matching scheme for 3D image volumes that supports finding an optimal threshold for detecting (near-) duplicates and non-duplicates. Our main contributions are as follows:
  • We introduce a matching approach for 3D image volumes that relies on the majority counts from slice-wise retrievals, weakening the impacts of background slices on volume retrieval. We show how to identify an optimal threshold for both near-duplicate and duplicate image detection.
  • We synthesize near-duplicates using various image transformations with different distortion levels. The robustness of different pretrained feature embeddings towards these transformations is quantitatively benchmarked.
  • We evaluate the detection performance using embeddings extracted by DINO [10, 11], the state-of-the-art self-supervised models designed for natural images, demonstrating the transferability of such models to medical imaging tasks without fine-tuning as studied in [12].
II Methods
Our proposed method relies on retrieving 3D medical images based on slice-wise similarity scores and cross-volume aggregation. As shown in Figure 1, a 2D feature extractor $\phi$ is used to embed the 2D slices into a vector representation. Let $X^{(i)}$ be the image volume related to case $i$ in the database and $x^{(i)}_{j}:=\phi(X^{(i)}_{j})$ be the embedding of the $j$-th slice. In our implementation all embeddings are stored in a vector database using the FAISS backend [13]. Let $Q$ represent a query image volume consisting of $n$ slices, and $q_{m}:=\phi(Q_{m})$ be the embedding of the $m$-th query volume slice. For all query slices $q_{m}(m=1,...,n)$ we retrieve the embedding from the database which has minimal Euclidean distance with $q_{m}$, and denote the corresponding case with $N[q_{m}]$. A histogram is built for all $n$ query slices to identify the case ID in the database with the most hits:
$h[i]=\sum_{m=1}^{n}\mathds{I}(N[q_{m}],X^{(i)})$ (1)
Here, $\mathds{I}$ is the identity function that returns 1 if $N[q_{m}]$ matches $X^{(i)}$ and 0 otherwise. The normalized count for top-$k$ predictions is then defined as
$c(k)=\frac{\sum_{l=1}^{k}h[i_{l}]}{n}$ (2)
where $i_{1},...,i_{k}$ denote the indices of the $k$ largest hits, i.e. $h[i_{1}]\geq...\geq h[i_{k}]\geq h[j]$ for all $j\not\in\{i_{1},...,i_{k}\}$. The normalized count has a range between 0 and 1. Based on $c(k)$ and a threshold, (near-) duplicates and non-duplicates will be identified. We conduct experiments to finetune this threshold for optimal detection performance and propose in Algorithm 1 a method to select an optimal threshold $t_{\textrm{opt}}$ across different query sets $\mathcal{Q}:=\left\{s_{1},s_{2},...,s_{N}\right\}$ which are defined in Section IV-A3. First, we generate a candidate list $\mathbf{T}$ of thresholds based on the maximal Youden’s index associated with the Receiver Operating Characteristic (ROC) curve of each query set $s\in\mathcal{Q}$. The final optimal threshold $t_{\textrm{opt}}$ is then chosen to maximize the sum of average sensitivity and specificity across all query sets.
Algorithm 1 Finding optimal decision threshold
1:$\mathcal{Q}:=\left\{s_{1},s_{2},...,s_{N}\right\}$
2:init array $\mathbf{T}$, $\mathbf{SE}$, $\mathbf{SP}$
3:for $u=1,2,...,N$ do
4:     $R_{u}:=$ ROC curve for $s_{u}$
5:     $\mathbf{T}[u]\leftarrow t_{ROC}:=\underset{r\in R_{u}}{\textrm{argmax}}(\textrm{sens}[r]+\textrm{spec}[r])$
6:     for $v=1,2,...,N$ do
7:         $\mathbf{SE}[u,v]\leftarrow$ sensitivity of $s_{v}$ at threshold $\mathbf{T}[u]$
8:         $\mathbf{SP}[u,v]\leftarrow$ specificity of $s_{v}$ at threshold $\mathbf{T}[u]$
9:     end for
10:end for
11:$t_{\textrm{opt}}\leftarrow\mathbf{T}\left[\underset{u=1,...,N}{\textrm{argmax}}\frac{1}{N}\sum_{v=1}^{N}\left(\mathbf{SE}[u,v]+\mathbf{SP}[u,v]\right)\right]$
Figure 1: Retrieval at case level based on count accumulation
III Experiments
III-A Datasets
The Medical Segmentation Decathlon (MSD) [14] is a biomedical image analysis challenge comprising segmentation tasks related to different body organs and imaging modalities. The images are either CT or MRI of different sequences that were acquired across multiple institutions and represent real-world data. The dataset was originally split by the challenge organizer into a train and test set which consist of 1723 and 887 cases respectively. To build our benchmark, we further split the MSD train set into 3 buckets 1A, 1B, 1C, and the MSD test set into 3 buckets 2A, 2B, 2C as indicated in Table I. Buckets 1A/2A consist of the first 50% of train/test cases across all tasks and are used to build the database and query for duplicate images. Buckets 1B/2B contain the same cases as buckets 1A/2A but with different image transformations to generate near-duplicate images. Buckets 1C/2C consist of the second 50% of train/test cases and serve as the query for non-duplicate images. We considered each case as a unique patient and leveraged embeddings of different feature extractors, see Section III-B, to retrieve similar images of the query patient in the database. We used buckets 1A/1B/1C to find the decision threshold for (near-) duplicate and non-duplicate classification (see Algorithm 1) and evaluated this threshold on bucket 2A/2B/2C (Table V). Buckets 1B and 2B contain synthetic near-duplicate images by applying the following transformations with package Scipy on original images with different strengths:
  • Cropping: crop border in x, y, z dimensions with 5%, 10%, 15%, and 20% of the original full size
  • Rotation: rotate the (x, y) plane by 5, 10, 15, and 20 degrees
  • Translation: shift the images in x and y dimensions by 5%, 10%, 15%, 20% of the original size
  • Blurring: blur images with increasing sigma of 1, 2, 4, and 8
  • JPEG Compression: compress images under different quality of 100%, 75%, 50%, and 25%
  • Gaussian noise: add Gaussian noise with increasing standard deviation of 0.1, 0.2, 0.4 and 0.8
The first level of strength across all transformations aims to simulate realistic preprocessing steps applied to datasets while the rest 3 levels serve as extreme cases to test the robustness of feature embeddings. Figure 2 shows examples of transformed heart MRI images.
TABLE I: The number of cases per bucket
Bucket 1 Bucket 2 Usage
A 859 443 Database & duplicate query
B 859 443 Synthetic near-duplicate query
C 864 444 Non-duplicate query
Figure 2: Near-duplicate images generated by geometric and intensity transformations under different strengths. For JPEG compression, a region of interest is shown to display the quality degradation which is not visible under human eyes when looking at the whole image.
III-B Feature extraction and indexing
Our selected feature extractors are DINOv1 [10] and DINOv2 [11], which are powerful self-supervised pretrained models on natural images. The image preprocessing steps included min-max scaling of images to 0-1 and resizing the axial slices to 224$\times$224. Embeddings in buckets 1A and 2A are indexed with Locality-Sensitivity Hashing (LSH) [15] and Hierarchical Navigable Small World (HNSW) [16] for effective similarity search.
III-C Evaluation
We evaluated the classification in two stages. In the first stage, a query image is considered duplicate if $c(k)$ surpasses a threshold value $t$. Here we considered $c(1)$ and $c(3)$ for top-1 and top-3 retrievals, respectively. We used the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC) to evaluate the class separability at different threshold values. In the second stage, we took the threshold $t_{ROC}$, defined in line 5 of Algorithm 1, to classify the queries into positive and negative outputs. Subsequently, we compared the highest-count case ID $X^{(i_{1})}$ from each case classified as duplicate to the ground-truth case IDs. If a match is found, the query output is classified as a true positive. Otherwise, it is labeled as a false positive. We computed in this stage the sensitivity and specificity values associated with $t_{ROC}$.
IV Results
IV-A Decision threshold determination
In this section, we analyzed embeddings from buckets 1A, 1B, and 1C to determine the best threshold for (near-) duplicate detection using Algorithm 1.
IV-A1 Duplicate vs. Non-duplicate
As an overview of all thresholds, Table II displays the AUC for the first stage evaluation. The top-1 retrieval achieves AUC in the range of 0.98-0.99, while the top-3 retrievals exhibit a slight decrease to 0.96-0.98. Figure 3 illustrates the accumulated count distribution for top-1 and top-3 retrievals indexed by HNSW using DINOv1 embeddings. It is evident that extending from top-1 to top-3 retrievals results in a significant increase in counts, boosting sensitivity while decreasing specificity due to higher false positives. We identified $t_{ROC}$ for all top-k retrievals and computed the associated sensitivity and specificity for the second stage (Table II). Stage 2 maintains a stable sensitivity rate of 0.99 across top-k retrievals, with slightly lower specificity in the top 3 due to increased false positives in the first stage. However, false positives stemming from case ID mismatches in the second stage are relatively low. We also observed that HNSW indexing offers slightly higher sensitivity than LSH, while LSH provides higher specificity compared to HNSW. In terms of extractors, DINOv2 outperformed DINOv1 by a small margin.
TABLE II: (Bucket 1A-1C) Duplicate detection evaluation
DINOv1 DINOv2
LSH HNSW LSH HNSW
AUC Top 1 0.9884 0.9873 0.9901 0.9898
(Stage 1) Top 3 0.9658 0.9834 0.9681 0.9888
Sensitivity Top 1 0.9930 0.9942 0.9965 0.9953
(Stage 2) Top 3 0.9965 0.9988 0.9977 0.9988
Specificity Top 1 0.9295 0.9225 0.9353 0.9318
(Stage 2) Top 3 0.9306 0.9156 0.9353 0.9272
Figure 3: Case-level normalized counts of top 1 and top 3 predictions indexed with HNSW using DINOv1 embeddings. Green bars denote non-duplicate queries and gray and red bars denote duplicate queries. Red bars show duplicate queries in which the top 1 predictions are in the database but do not match the ground truth image case ID. The blue horizontal bar is the threshold resulting in the maximum sum of sensitivity and specificity.
IV-A2 Near-duplicate vs. Non-duplicate
Table III displays the stage 1’s AUC of the top 3 retrievals indexed by HNSW and using DINOv1 embeddings. The ROC curves for top 1 with combinations of other indices and feature extractors follow the same tendency and hence are not shown. Under lowest strength, cropped, translated, noisy, and JPEG-compressed embeddings maintained AUC scores above 0.95, followed by rotation at 0.9327 and blurring at 0.9360. When stronger transformations were considered, all transformations showed an expected decrease in AUC scores proportional to the strength. At the strongest level, the AUC scores by blurring and Gaussian noise rose by a small margin compared to the preceded level. We postulated that under this range, the images are significantly altered and the embeddings are matched to non-informative slices of different volumes. Therefore even though the AUC scores seem to increase in stage 1, they show a consistent decrease in stage 2 where case IDs are matched (Table IV). Regarding the feature extractors, we observed a lower performance of DINOv2 compared to DINOv1 by a margin of 0.1-0.3 AUC. Nevertheless, the best and worst transformations are consistent between the two types of embeddings.
TABLE III: (Bucket 1B-1C, Stage 1) AUC of near-duplicates classification using DINOv1 embeddings with HNSW index for top 3 retrieval
Transform Strength of transformation
Cropping crop=5% crop=10% crop=15% crop=20%
0.9830 0.9618 0.8817 0.8024
Rotation angle=5o angle=10o angle=15o angle=20o
0.9327 0.9053 0.8645 0.8426
Translation shift=5% shift=10% shift=15% shift=20%
0.9619 0.9493 0.9388 0.9209
Blurring sigma=1 sigma=2 sigma=4 sigma=8
0.9360 0.7127 0.6916 0.7711
JPEG quality=100% quality=75% quality=50% quality=25%
compression 0.9821 0.9699 0.9405 0.8816
Gaussian std=0.1 std=0.2 std=0.4 std=0.8
noise 0.9675 0.8437 0.6456 0.6789
TABLE IV: (Bucket 1B-1C, Stage 2) Sensitivity of near-duplicates classification using DINOv1 embeddings with HNSW index for top 3 retrieval
Transform Strength of transformation
Cropping crop=5% crop=10% crop=15% crop=20%
0.9942 0.9604 0.8161 0.6286
Rotation angle=5o angle=10o angle=15o angle=20o
0.8778 0.8522 0.8300 0.8114
Translation shift=5% shift=10% shift=15% shift=20%
0.9581 0.8708 0.8545 0.8510
Blurring sigma=1 sigma=2 sigma=4 sigma=8
0.8196 0.5239 0.1735 0.0163
JPEG quality=100% quality=75% quality=50% quality=25%
compression 0.9709 0.9267 0.8468 0.8102
Gaussian std=0.1 std=0.2 std=0.4 std=0.8
noise 0.9779 0.8568 0.3050 0.0664
IV-A3 Selected optimal threshold
After observing the retrieval performance using buckets 1A, 1B, and 1C, we decided to prioritize the top 3 retrievals using the HNSW index. Even though this choice slightly decreases specificity, it offers higher sensitivity as a trade-off. We determined $t_{opt}$ on the train set (Algorithm 1) with 7 query sets, including 6 near-duplicate query sets of the lowest transformation strength and 1 duplicate query set. The chosen $t_{opt}$ is 0.7711 and will be used to evaluate the detection performance using buckets 2A (duplicates), 2B (near-duplicates), and 2C (non-duplicates).
IV-B Evaluation with selected decision threshold
In Table V, we present the best sensitivity and specificity results for (near-) duplicate detection on buckets 2A, 2B, 2C using the top-3 retrievals obtained through DINOv1 embeddings with HNSW indexing. Overall, for stage 1, the mean sensitivity score is 0.9645, while for stage 2, it is 0.9407. In terms of mean specificity, we found scores of 0.8559 for stage 1 and 0.8373 for stage 2. Rotation exhibits the lowest sensitivity, mirroring trends observed in the training dataset. We also observed some false positive occurrences that are indeed in our database, which is due to the fact that possible identical images are shared among different MSD tasks. Upon scanning through 10 tasks, we were able to find 28 pairs of possible (near-) duplicates in the original MSD dataset shown in Table B.1. The majority of detected (near-) duplicates are from task 1, which contains the MRI brain images. An example is shown in Figure B.1 where two cases are different only in the brightness.
TABLE V: (Bucket 2A-2B-2C) Sensitivity and specificity using the threshold of 0.7711 for DINOv1 embeddings with HNSW index
Sensitivity Specificity
Transform Stage 1 Stage 2 Stage 1 Stage 2
No transform (duplicate) 0.9977 0.9977 0.8559 0.8559
Cropping (crop=5%) 0.9865 0.9865 0.8559 0.8559
Translation (shift=5%) 0.9481 0.9481 0.8559 0.8559
Rotation (angle=5o) 0.8849 0.8826 0.8559 0.8539
Blurring (sigma=1) 0.9639 0.8442 0.8559 0.7646
JPEG (quality=100%) 0.9977 0.9549 0.8559 0.8207
Noise (std=0.1) 0.9729 0.9707 0.8559 0.8539
Mean 0.9645 0.9407 0.8559 0.8373
V Conclusion
In this study, we introduced and evaluated an unsupervised approach for near- and duplicate detection of 3D medical images relying on pretrained 2D vision embeddings. We quantitatively assessed its performance to detect near-duplicates involving various image transformations and different strength levels. A heuristic based on Youden’s index is proposed to select a single threshold value for balanced performance in the (near-) duplicate detection task. Our top results show 0.9645 sensitivity and 0.8559 specificity for detecting duplicates and near-duplicates. This underscores the transferability of pretrained embeddings from natural images to medical imaging. Additionally, our method successfully pinpointed potential (near-) duplicates in MSD datasets, validating its effectiveness in identifying real-world duplicates. We consider extending our study with in-depth analysis under different datasets, modalities, contrasts, and organs, as well as investigating naturally occurring near duplicates, e.g., same patient at different time points, together with different volume sampling strategies to address spatial redundancy in future work.
Appendix
V-A Near-duplicate detection results
We present in this section the full results of near-duplicate analysis on buckets 1A, 1B, and 1C. The evaluation was conducted using two feature extractors DINOv1 and DINOv2, and two database indices LSH and HNSW. For stage 1 (Table A.1), we report only AUC, and for stage 2, we report the sensitivity (Table A.2) and specificity (Table A.3) associated with $t_{ROC}$. For all metrics, we show the results with top-1 and top-3 retrieval.
TABLE A.1: (Bucket 1B-1C, Stage 1) AUC of near-duplicates classification using DINOv1 and DINOv2 embeddings with LSH and HNSW indexing. Left: top-1 retrieval. Right: top-3 retrieval
Transform Strength DINOv1 DINOv2
LSH HNSW LSH HNSW
Crop crop=5% 0.9902 0.9887 0.9661 0.9732
crop=10% 0.9751 0.9762 0.8964 0.9102
crop=15% 0.8736 0.8932 0.7624 0.7835
crop=20% 0.7876 0.7989 0.7190 0.7193
Rotation angle=5o 0.9428 0.9457 0.9421 0.9461
angle=10o 0.9156 0.9202 0.8864 0.8864
angle=15o 0.8791 0.8775 0.8092 0.7912
angle=20o 0.8576 0.8621 0.7225 0.6874
Translation shift=5% 0.9688 0.9678 0.9675 0.9672
shift=10% 0.9578 0.9577 0.9587 0.9580
shift=15% 0.9474 0.9476 0.9482 0.9447
shift=20% 0.9270 0.9290 0.9327 0.9273
Blurring sigma=1 0.9385 0.9437 0.9008 0.8940
sigma=2 0.7031 0.7164 0.6774 0.6529
sigma=4 0.6330 0.6778 0.7456 0.6848
sigma=8 0.7043 0.7534 0.8232 0.7470
quality=100% 0.9820 0.9810 0.9830 0.9811
JPEG quality=75% 0.9736 0.9741 0.9636 0.9618
Compression quality=50% 0.9480 0.9504 0.9344 0.9247
quality=25% 0.8745 0.8894 0.8422 0.8562
std=0.1 0.9621 0.9681 0.9292 0.9380
Gaussian std=0.2 0.8195 0.8485 0.7956 0.8020
noise std=0.4 0.6069 0.6354 0.5976 0.5951
std=0.8 0.6451 0.6571 0.6114 0.6248
Transform Strength DINOv1 DINOv2
LSH HNSW LSH HNSW
Crop crop=5% 0.9579 0.9830 0.9351 0.9581
crop=10% 0.9372 0.9618 0.8661 0.8807
crop=15% 0.8528 0.8817 0.7622 0.7781
crop=20% 0.7906 0.8024 0.7276 0.7282
Rotation angle=5o 0.9579 0.9830 0.9351 0.9581
angle=10o 0.9372 0.9618 0.8661 0.8807
angle=15o 0.8528 0.8817 0.7622 0.7781
angle=20o 0.7906 0.8024 0.7276 0.7282
Translation shift=5% 0.9411 0.9619 0.9399 0.9593
shift=10% 0.9278 0.9493 0.9260 0.9448
shift=15% 0.9185 0.9388 0.9128 0.9293
shift=20% 0.9013 0.9209 0.8970 0.9124
Blurring sigma=1 0.9164 0.9360 0.8811 0.8817
sigma=2 0.7005 0.7127 0.7013 0.6773
sigma=4 0.6561 0.6916 0.7752 0.7267
sigma=8 0.7191 0.7711 0.8515 0.7902
quality=100% 0.9637 0.9821 0.9639 0.9841
JPEG quality=75% 0.9474 0.9699 0.9385 0.9535
Compression quality=50% 0.9185 0.9405 0.9038 0.9141
quality=25% 0.8531 0.8816 0.8213 0.8328
std=0.1 0.9329 0.9675 0.9015 0.9090
Gaussian std=0.2 0.8002 0.8437 0.7696 0.7618
noise std=0.4 0.6217 0.6456 0.6104 0.5814
std=0.8 0.6573 0.6789 0.6296 0.6190
TABLE A.2: (Bucket 1B-1C, Stage 2) Sensitivity of near-duplicates classification using DINOv1 and DINOv2 embeddings with LSH and HNSW indexing. Left: top-1 retrieval. Right: top-3 retrieval
Transform Strength DINOv1 DINOv2
LSH HNSW LSH HNSW
Crop crop=5% 0.9732 0.9208 0.9627 0.6321
crop=10% 0.9464 0.9546 0.6764 0.7229
crop=15% 0.7695 0.7637 0.4517 0.4319
crop=20% 0.4843 0.5204 0.3504 0.3667
Rotation angle=5o 0.8952 0.9371 0.9313 0.9348
angle=10o 0.8359 0.8044 0.7416 0.5460
angle=15o 0.7811 0.7765 0.2992 0.2794
angle=20o 0.7625 0.7590 0.1746 0.1513
Translation shift=5% 0.9488 0.9756 0.9418 0.9686
shift=10% 0.9034 0.9395 0.8917 0.9139
shift=15% 0.8685 0.9127 0.8417 0.8533
shift=20% 0.8405 0.8591 0.8231 0.8300
Blurring sigma=1 0.7998 0.7974 0.7509 0.7683
sigma=2 0.3760 0.4144 0.1735 0.1921
sigma=4 0.0861 0.3539 0.0442 0.0373
sigma=8 0.0093 0.0035 0.0116 0.0023
quality=100% 0.9383 0.9464 0.9302 0.9360
JPEG quality=75% 0.9302 0.8941 0.8766 0.8091
Compression quality=50% 0.8778 0.8941 0.8172 0.8149
quality=25% 0.7870 0.8056 0.6973 0.7055
std=0.1 0.9616 0.9325 0.8917 0.7322
Gaussian std=0.2 0.7113 0.7474 0.6228 0.6403
noise std=0.4 0.2701 0.2864 0.2375 0.2177
std=0.8 0.0268 0.0326 0.0384 0.0349
Transform Strength DINOv1 DINOv2
LSH HNSW LSH HNSW
Crop crop=5% 0.9942 0.9942 0.7322 0.7288
crop=10% 0.9325 0.9604 0.6310 0.7159
crop=15% 0.7800 0.8161 0.3935 0.3970
crop=20% 0.5972 0.6286 0.3411 0.3283
Rotation angle=5o 0.8731 0.8778 0.8417 0.8359
angle=10o 0.8428 0.8522 0.5274 0.5320
angle=15o 0.8161 0.8300 0.3143 0.3399
angle=20o 0.8068 0.8114 0.2456 0.2212
Translation shift=5% 0.9430 0.9581 0.8987 0.8941
shift=10% 0.9010 0.8708 0.8568 0.8312
shift=15% 0.8487 0.8545 0.7497 0.7718
shift=20% 0.8428 0.8510 0.7485 0.8126
Blurring sigma=1 0.8091 0.8196 0.6938 0.6868
sigma=2 0.5704 0.5239 0.3120 0.2305
sigma=4 0.1538 0.1735 0.0675 0.0477
sigma=8 0.0256 0.0163 0.0221 0.0128
quality=100% 0.9686 0.9709 0.9197 0.9616
JPEG quality=75% 0.9499 0.9267 0.8789 0.8533
Compression quality=50% 0.8440 0.8498 0.6985 0.6612
quality=25% 0.7858 0.8102 0.5844 0.5390
std=0.1 0.9837 0.9779 0.7811 0.4924
Gaussian std=0.2 0.8207 0.8568 0.6286 0.4889
noise std=0.4 0.2992 0.3050 0.2398 0.1665
std=0.8 0.0640 0.0664 0.0629 0.0419
TABLE A.3: (Bucket 1B-1C, Stage 2) Specificity of near-duplicates classification using DINOv1 and DINOv2 embeddings with LSH and HNSW indexing. Left: top-1 retrieval. Right: top-3 retrieval.
Transform Strength DINOv1 DINOv2
LSH HNSW LSH HNSW
Crop crop=5% 0.8679 0.9734 0.8885 0.9746
crop=10% 0.8602 0.8467 0.8848 0.8693
crop=15% 0.7659 0.8344 0.7991 0.8731
crop=20% 0.4805 0.5306 0.6083 0.6434
Rotation angle=5o 0.8061 0.7768 0.8352 0.8185
angle=10o 0.8231 0.8925 0.8609 0.9109
angle=15o 0.8754 0.8776 0.8851 0.8963
angle=20o 0.8385 0.8441 0.8503 0.8534
Translation shift=5% 0.8667 0.8345 0.8883 0.8603
shift=10% 0.8576 0.8343 0.8730 0.8554
shift=15% 0.8567 0.8364 0.8730 0.8534
shift=20% 0.8557 0.8307 0.8720 0.8573
Blurring sigma=1 0.7265 0.7072 0.8503 0.8486
sigma=2 0.6660 0.6211 0.7151 0.6647
sigma=4 0.3539 0.2916 0.4315 0.4114
sigma=8 0.3241 0.3453 0.3725 0.4103
quality=100% 0.8626 0.8344 0.8846 0.8613
JPEG quality=75% 0.8525 0.8931 0.8814 0.9091
Compression quality=50% 0.8348 0.8077 0.8642 0.8429
quality=25% 0.7757 0.7598 0.8203 0.8190
std=0.1 0.8534 0.9087 0.8834 0.9168
Gaussian std=0.2 0.6831 0.6639 0.7895 0.7695
noise std=0.4 0.2459 0.3225 0.4099 0.4854
std=0.8 0.3956 0.3979 0.5412 0.5574
Transform Strength DINOv1 DINOv2
LSH HNSW LSH HNSW
Crop crop=5% 0.9133 0.9133 0.9260 0.9272
crop=10% 0.8694 0.8451 0.8798 0.8707
crop=15% 0.8326 0.8466 0.8527 0.8711
crop=20% 0.5804 0.6254 0.7112 0.7180
Rotation angle=5o 0.8705 0.8590 0.8798 0.8890
angle=10o 0.8555 0.8497 0.8738 0.8773
angle=15o 0.8197 0.7968 0.8566 0.8472
angle=20o 0.7569 0.7344 0.8121 0.7968
Translation shift=5% 0.8705 0.8532 0.8809 0.8809
shift=10% 0.8705 0.9017 0.8798 0.9121
shift=15% 0.9168 0.9133 0.9295 0.9272
shift=20% 0.9179 0.9110 0.9306 0.9249
Blurring sigma=1 0.7351 0.7330 0.8611 0.8661
sigma=2 0.4257 0.4990 0.5181 0.5772
sigma=4 0.2776 0.3527 0.3580 0.4696
sigma=8 0.2775 0.3183 0.3502 0.3959
quality=100% 0.9211 0.9000 0.9256 0.9188
JPEG quality=75% 0.8459 0.8740 0.8609 0.8915
Compression quality=50% 0.8840 0.8819 0.8971 0.8982
quality=25% 0.8083 0.8090 0.8412 0.8393
std=0.1 0.8705 0.8971 0.8798 0.9109
Gaussian std=0.2 0.6626 0.7050 0.7762 0.7910
noise std=0.4 0.3152 0.3764 0.5063 0.5726
std=0.8 0.4276 0.4423 0.5961 0.6657
V-B Detected (near-) duplicates in MSD dataset
We scanned through the MSD datasets using our proposed approach to detect near- and duplicates across tasks and train/test subsets. We used the DINOv1 embeddings indexed with HNSW and a threshold of 0.8. Any potential (near-) duplicate flagged by the algorithm was manually reviewed and confirmed. We presented in the tables below the case IDs that were flagged as (near-) duplicates and the subset to which they belong. Case ID 1 and Case ID 2 refer to the filenames of duplicate pairs in the dataset.
TABLE B.1: Detected (near-) duplicates in MSD dataset
Case ID 1 Subset Case ID 2 Subset
BRATS_432 Train BRATS_142 Train
BRATS_416 Train BRATS_099 Train
BRATS_442 Train BRATS_166 Train
BRATS_398 Train BRATS_063 Train
BRATS_082 Train BRATS_404 Train
BRATS_460 Train BRATS_225 Train
BRATS_065 Train BRATS_400 Train
BRATS_406 Train BRATS_084 Train
BRATS_445 Train BRATS_172 Train
BRATS_464 Train BRATS_235 Train
BRATS_396 Train BRATS_051 Train
BRATS_224 Train BRATS_459 Train
BRATS_457 Train BRATS_220 Train
BRATS_405 Train BRATS_083 Train
BRATS_448 Train BRATS_181 Train
BRATS_261 Train BRATS_471 Train
BRATS_550 Train BRATS_549 Train
BRATS_666 Train BRATS_663 Train
BRATS_508 Train BRATS_509 Train
BRATS_522 Train BRATS_521 Train
BRATS_717 Test BRATS_226 Train
BRATS_691 Test BRATS_052 Train
BRATS_720 Test BRATS_243 Train
BRATS_710 Test BRATS_169 Train
BRATS_719 Test BRATS_233 Train
BRATS_685 Test BRATS_030 Train
BRATS_721 Test BRATS_253 Train
BRATS_687 Test BRATS_032 Train
BRATS_727 Test BRATS_270 Train
BRATS_683 Test BRATS_022 Train
BRATS_701 Test BRATS_130 Train
BRATS_690 Test BRATS_047 Train
BRATS_715 Test BRATS_221 Train
BRATS_726 Test BRATS_268 Train
BRATS_702 Test BRATS_135 Train
BRATS_725 Test BRATS_262 Train
BRATS_711 Test BRATS_170 Train
BRATS_689 Test BRATS_042 Train
BRATS_693 Test BRATS_077 Train
BRATS_698 Test BRATS_117 Train
BRATS_684 Test BRATS_025 Train
BRATS_712 Test BRATS_194 Train
BRATS_724 Test BRATS_259 Train
BRATS_716 Test BRATS_223 Train
Case ID 1 Subset Case ID 2 Subset
BRATS_676 Test BRATS_021 Train
BRATS_707 Test BRATS_153 Train
BRATS_697 Test BRATS_116 Train
BRATS_728 Test BRATS_272 Train
BRATS_677 Test BRATS_024 Train
BRATS_699 Test BRATS_126 Train
BRATS_713 Test BRATS_198 Train
BRATS_723 Test BRATS_257 Train
BRATS_709 Test BRATS_163 Train
BRATS_678 Test BRATS_023 Train
BRATS_688 Test BRATS_034 Train
BRATS_708 Test BRATS_161 Train
BRATS_682 Test BRATS_027 Train
BRATS_694 Test BRATS_092 Train
BRATS_686 Test BRATS_031 Train
BRATS_695 Test BRATS_112 Train
BRATS_700 Test BRATS_128 Train
BRATS_704 Test BRATS_144 Train
BRATS_681 Test BRATS_028 Train
BRATS_714 Test BRATS_218 Train
BRATS_722 Test BRATS_254 Train
BRATS_692 Test BRATS_076 Train
BRATS_705 Test BRATS_146 Train
BRATS_703 Test BRATS_143 Train
BRATS_679 Test BRATS_029 Train
BRATS_696 Test BRATS_115 Train
BRATS_718 Test BRATS_230 Train
liver_14 Train liver_15 Train
liver_0 Train liver_1 Train
liver_171 Train liver_172 Train
liver_169 Train liver_170 Train
liver_137 Test liver_74 Train
hepaticvessel_265 Train spleen_19 Train
hepaticvessel_286 Train hepaticvessel_287 Train
hepaticvessel_376 Test spleen_18 Train
hepaticvessel_045 Test hepaticvessel_139 Train
spleen_1 Test hepaticvessel_433 Train
spleen_32 Train colon_041 Train
spleen_48 Train colon_055 Train
spleen_7 Test hepaticvessel_362 Train
pancreas_027 Test pancreas_032 Train
pancreas_007 Test pancreas_228 Train
colon_067 Test spleen_59 Train
colon_071 Test spleen_60 Train
Figure B.1: Example of a near-duplicate found in MSD dataset. The BRATS_404 and BRATS_082 are a near-duplicate pair that is different only in the brightness.